Geoffrey is twice as old as his friend Peter. Peter is 5 years older than Alice. In 5 years, Geoffrey will be three times as old as Alice. How old is Peter now?
Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Geoffrey is twice as old as his friend Peter. Peter is 5 years older than Alice. In 5 years, Geoffrey will be three times as old as Alice. How old is Peter now?
-----
Equations:
g = 2P
P = A+5
g+5 = 3(A+5)
--------
Substitute for "g" and "A" to get::
2P +5 = 3(P-5)+3*5
2P + 5 = 3P
P = 5 yrs (Peters age now)
-----------------
Cheers,
Stan H.
---------------

Answer by ikleyn(52884)

.
Geoffrey is twice as old as his friend Peter. Peter is 5 years older than Alice.
In 5 years, Geoffrey will be three times as old as Alice. How old is Peter now?
~~~~~~~~~~~~~~~~~~~

G = 2P              (1)
P = A + 5           (2)
G+5 = 3*(A+5)       (3)


===================> Substitute (1) into equation (3). In this way you eliminate unknown G 
                     and obtain the modified system for 2 unknowns A and P:


P = A + 5           (2')
2P + 5 = 3(A+5)     (3')


Now substitute (2') into (3'). You will get

2*(A+5) + 5 = 3(A+5)  ====>  2A + 15 = 3A + 15 ====>  A = 0.

Thus Alice is a newborn.

Then  from (2)  P = A + 5 = 0+5 = 5.


Answer. Peter is 5 years old.